Missing no. - 4,18,?,100,180,294
Answer by Edwin McCravy(20081)

  4 = 1×2²
 18 = 2×3²
  ?  
100 = 4×5²
180 = 5×6²   
294 = 6×7²

So obviously, to fill in the missing number:

  4 = 1×2²
 18 = 2×3²
 48 = 3×4²
100 = 4×5²
180 = 5×6²   
294 = 6×7²
100 = 4×5²
180 = 5×6²   
294 = 6×7²

Edwin